For the given function f and g, find the indicated composition.
f(x)=-5x+6, g(x)=2x+5
(g.f)(x)

Answer by jim_thompson5910(35256)

For the given functions f and g , find the indicated composition.
f(x) = -5x + 6, g(x) = 2x + 5
(g.f)(x)
--------------------------------------------------------------------------------------------------------------

Assuming you mean (g * f)(x)

(g * f)(x) is the same as g(x)*f(x)

So we multiply the functions to get

	g(x)*f(x) = (2x + 5)*(-5x + 6)
	
	
	g(x)*f(x) = 2x*(-5x + 6) + 5*(-5x + 6)
	
	
	g(x)*f(x) = 2x*(-5x) + 2x*(6) + 5*(-5x) + 5*(6)
	
	
	g(x)*f(x) = -10x^2 + 12x - 25x + 30

	
	g(x)*f(x) = -10x^2 - 13x + 30
	
	
	(g * f)(x) = -10x^2 - 13x + 30


The final answer is (g * f)(x) = -10x^2 - 13x + 30